find the slope and y-intercept fro each line tha thas a slope and y-intercept:
x+2y=3
thank you

Answer by Alan3354(69443)

find the slope and y-intercept fro each line tha thas a slope and y-intercept:
x+2y=3
---------
Put into slope-intercept form, y = mx + b (m = slope, b = y-intercept)
2y = -x + 3
y = (-1/2)x + 3/2
Slope = -1/2, y-intercept = 3/2